OpenStudy (anonymous):

The transformation of the function f(x)=(1/2)^x when it becomes f(x)=(1/2)^x+3 is a horizontal shift of 3 units to the right. True or false?

OpenStudy (anonymous):

false

OpenStudy (fifciol):

false, this would be true f(x)=(1/2 )^(x-3)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

@ashlyn17 You might wanna keep this in mind while solving transformation questions: If original function is Y=f(x): Y=f(x)+a implies vertical shift a units Y=f(x+a) horizontal shift -a units f(ax) horizontal shrink (or stretch if |a|<1) by a units (divide the x's by a) a•f(x) vertical stretch (or shrink if |a|<1) by a factor of a f(-x) reflect over the y-axis -f(x) reflect over the x-axis